A computer sends a packet of information along a channel and waits for a return signal acknowledging that the packet has been received. If no acknowledgment is received within a certain time, the packet is re-sent. Let X represent the number of times the packet is sent. Assume that the probability mass function of X is given by
p(x) ={ cx , for x=0,1,2,...5
0 , otherwie
Find P(X = 2).

Answer by rothauserc(4718)

x = 0, 1, 2, 3, 4, 5
:
p(x) for x = 0, 1, 2, 3, 4, 5 = 0, c, 2c, 3c , 4c, 5c = 1
:
0 + c + 2c + 3c + 4c + 5c = 1
:
15c = 1
:
c = 1/15
:
p(x) for x = 0, 1, 2, 3, 4, 5 = 0, 1/15, 2/15, 3/15, 4/15, 5/15
:
***************
P(X = 2) = 2/15
***************
: